A student studying the photoelectric effect for two different metals records the following information.

(i) the stopping potential for photoelectrons released from metal 1 is 1.24 V larger than that for metal 2
(ii) the threshold frequency for metal 1 is 50.0% smaller than that for metal 2.

Determine the work function for each metal.

ϕ1 =

ϕ2 =
Answer #1


here V01 = -1.24 V02

V01 = -0.5 v02

eV0 = hf - hvo

hvp = evo1 + hvo1 for metal 1

hvo = evo2 + hv02 for metal 2


from frist two equations

evo1 + hv01 = evo2 + hvo2

e(1.24 v02) + h(0.5 v02) = ev02 + hv02   

e(0.24 vo2) + h(0.5 vo2)

hv02 = 24/50 ev02

phi = hv0 = 24/50 ev02

phi1 = 0.50 * 24/50 = 0.24 * 1.6 e -19 V02 = 0.384 Vo2

phi 2 = phi1/0.5 = 0.768 Vo2